OpenStudy (anonymous):

Factor completely: 3x2 + 9x − 3 3(x2 + 3) 3(x2 + 3x − 1) 3x(x2 + 3x − 1) Prime

OpenStudy (anonymous):

Somebody please help.

OpenStudy (anonymous):

3(x^2+3x-1)

OpenStudy (anonymous):

theres 2 x's with 3 so they get put inside parenthesis the lowest 9 can be broken down is 3x

OpenStudy (anonymous):

the 3 outside of the parenthesis gets left without an x

OpenStudy (anonymous):

\(\Large 3x^2+9x-3\\\Large =3x^2+3\times3x-3\times1\\\Large =3(x^2+3x-1)\)